# Runge-Kutta Approximations for Direct Coning Compensation Applying Lie Theory

### v1

@@ 3.1 Simplification of the Coning Correction
When the rotation vector is small (which it often is for small time steps since we always start the integration from $\boldsymbol{\phi}=\textbf{0}$), the coefficient of the right-hand term may be expanded as a Taylor series in the same was as eq. (15) to find
@@
which is an excellent approximation in most cases. More than this, we often make the so-called Goodman-Robinson approximation [16] that maintains only first-order terms in the attitude and replaces $\boldsymbol{\phi}$ with $\boldsymbol{\theta}$ inside the integrand
@@
The replacement of $\boldsymbol{\phi}$ by $\boldsymbol{\theta}$ may be viewed as a single iteration within a Picard iteration scheme. The Goodman-Robinson approximation, which predates the Bortz’s work by over a decade, is the foundation for many of the classical coning algorithms [1, 3, 2]. Some newer algorithms attempt to solve either eq. (24) or eq. (26), though this comes with an increase in complexity [17, 18, 19].
@@ 3.2 Approximate Solutions to the Coning Correction Term
Consider a rate integrating gyro producing measurements of $\Delta\boldsymbol{\theta}_{k}$ at the sensor interval from eq. (21). Suppose that $m$ such measurements $\{\Delta\boldsymbol{\theta}_{k}\}_{k=1}^{m}$ are accumulated over an IMU’s minor interval. The objective is to integrate this set of IMU observations to produce an estimate of the DCM $\emph{{T}}_{m}$ or the rotation vector $\boldsymbol{\phi}_{m}$ at the end of the minor interval. The Goodman-Robinson approximation from eq. (27) may be used to solve this problem analytically if we make some assumptions on the time-history of $\boldsymbol{\omega}$. The simplest non-trivial assumption is a linearly varying angular velocity.
Most of the classical coning literature focuses on the two-speed paradigm where $m>1$, where the “two speeds” being the rate of sensor interval and the rate of the minor interval (c.f. Fig. 1). Under the two-speed paradigm, where $m>1$ sensor interval measurements are obtained over a minor interval, we may assume a linear $\boldsymbol{\omega}$ to obtain the classic approximation [1]
@@
The two-speed techniques exemplified by eq. (28)—though there are many others under different assumptions for the time-history of $\boldsymbol{\omega}$ [3, 2]—are exceptionally fast and remove the need for performing the computationally expensive task of finite rotations at the sensor interval. However, there have been considerable advancements in computing since 1990, and two-speed algorithms are not always necessary. Propagating at the sensor interval is well-known to be desirable when it is possible [2].
If we choose to perform finite rotations at the sensor interval, then we are only concerned with the coning correction over the sensor interval (and not the minor interval). This leads to the one-speed algorithm, which may be easily derived for linear $\boldsymbol{\omega}$ as described in Appendix A and results in
@@
which is the same correction suggested by Miller in the early 1980s [20]. Thus, letting $\Delta\emph{{T}}_{k}=\emph{{T}}(\Delta\boldsymbol{\phi}_{k})$ be computed from eq. (4), the attitude at the end of the sensor interval is then computed as the product
$\emph{{T}}_{m}=\prod_{k=1}^{m}\Delta\emph{{T}}_{k}=\Delta\emph{{T}}_{m}\ldots\Delta\emph{{T}}_{k}\ldots\Delta\emph{{T}}_{1}$ (31)
@@ 4 Numerical Methods for Coning Correction
In light of eq. (19), coning compensation amounts to solving the differential equation of eq. (13) for $\boldsymbol{\phi}\left(t_{k}+\Delta t\right)$ from an initial condition $\boldsymbol{\phi}\left(t_{k}\right)=0$. This generalized perspective enables usage of traditional numerical solvers such as Runge-Kutta methods. Indeed, using traditional Runge-Kutta style solvers we can obtain identical expressions for the single-speed algorithm fom eq. (30). Our motivation is three-fold. First, equivalence with traditional coning correction methods demonstrates applicability of Runge-Kutta style solvers. Second, generalized solvers accommodate non-zero initial conditions and higher-order models without explicit derivations. Third, non-zero initial conditions are tolerated avoiding exponentiation as the sensor interval.
@@ 4.1 Runge-Kutta methods and sampling requirements
The Runge-Kutta algorithm, stated in Algorithm 1, is a generalized numerical solver for initial value problems. The ordinary differential equation $f(t,\boldsymbol{\phi}):\mathbb{R}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, is solved from an initial condition ${\boldsymbol{\phi}}_{k}\in\mathbb{R}^{3}$ at time $t_{k}$ to the final value ${\boldsymbol{\phi}}_{k+1}={\boldsymbol{\phi}}\left(t_{k}+\Delta t\right)$ at stepsize $\Delta t$.
@@
Algorithm 1 extends to different solver orders and sampling schemas as determined by $\emph{{A}}\in\mathbb{R}^{n\times n}$, $\emph{{b}},\emph{{c}}\in\mathbb{R}^{n}$. Butcher tableaux [21, 22] are useful for depicting the algorithm hyperparameters A,b,c with examples shown in Tables 1 - 4 for common implementations. Note, the array c, depicted in tableau’s first column, identifies both the solver order and the timesteps at which point the differential equation, $f$, is evaluated. The sample points are listed relative to step size $\Delta t$ (c.f. $t_{0}+\Delta t\emph{{c}}_{\nu}$} in Algorithm 1, line 5).
Table 1: Forward Euler Tableau
@@
Table 2: Explicit-Midpoint Tableau
@@
Table 3: Runge-Kutta Original Third Order Tableau
@@
Table 4: Runge-Kutta Original Fourth Order Tableau
@@
To solve eq. (19), using Algorithm 1, we employ eq. (13) as
$f(t,\boldsymbol{\phi})\coloneq\emph{{J}}^{-1}_{\boldsymbol{\phi}}\boldsymbol{\omega}(t).$ (32)
In this case, the $f$ separates as two terms. The first term, $\emph{{J}}^{-1}_{\boldsymbol{\phi}}$, depends on the dependent variable $\boldsymbol{\phi}$. The second term $\boldsymbol{\omega}(t)$, for our problem, is obtained through sensor measurements. Notice, Algorithm 1 samples $\boldsymbol{\omega}(t)$ at discrete points as determined by $t_{k}$, $\Delta t$, and c. We first expand Runge-Kutta methods assuming direct access to $\boldsymbol{\omega}$ for any $t$.
Applying the forward-Euler method, calling Algorithm 1 using Table 1, yields
@@
Motivated by eq. (19), we select $\boldsymbol{\phi}_{k}=0$ which effects $\emph{{J}}_{\boldsymbol{\phi}_{k}}^{-1}=I$}. Consequently,
@@
For higher-order methods, expanding Algorithm (1) becomes tedious. However, using the initial condition $\boldsymbol{\phi}(t_{k})=0$, the small-angle approximation eq. (16), and retaining only quadratic terms, we obtain
@@
Samples of $\boldsymbol{\omega}$ must be provided either through direct measurements (e.g. gyros providing instantaneous rates) or from a model (e.g. parameterized from integrated rate measurements).
@@ 4.2 Approximating $\boldsymbol{\omega}$ from $\Delta\boldsymbol{\theta}$
When $\boldsymbol{\omega}$ is available, eq. (40) may be used directly. However, for rate-integrating gyros, we must model $\boldsymbol{\omega}$ from $\Delta\boldsymbol{\theta}$ measurements. Consider a model for $\boldsymbol{\omega}$ that is polynomial in the elapsed time
$\boldsymbol{\omega}\left(t\right)=\sum_{i=1}^{Q}\emph{{p}}_{i}\left(t-t_{k}\right)^{i-1}.$ (41)
The model parameters $\emph{{P}}_{Q}\in\mathbb{R}^{3\times Q}$
$\emph{{P}}_{Q}=\begin{bmatrix}\emph{{p}}_{1}&\emph{{p}}_{2}&\cdots&\emph{{p}}_{Q}\end{bmatrix}$ (42)
can be obtained from $\Delta\boldsymbol{\theta}$ measurements by solving a linear system of equations. For each measurement, we obtain an equation by replacing $\boldsymbol{\omega}$ in eq. (21) with eq. (41)
@@
$=\sum_{i=0}^{Q}\emph{{p}}_{k}\frac{1}{i}\left(\left(j\Delta t\right)^{i}-\left(j-1\right)^{i}\Delta t^{i}\right).$ (43b)
From two measurements, we can form an affine model for $\boldsymbol{\omega}$. Collecting two measurements in the matrix $\Theta_{2}\in\mathbb{R}^{3\times 2}$
$\Theta_{2}=\begin{bmatrix}\Delta\boldsymbol{\theta}_{k-1}&\Delta\boldsymbol{\theta}_{k}\end{bmatrix}$, (44)
we form a linear system of equations
$\Theta^{T}=\begin{bmatrix}-\frac{1}{2}\Delta t^{2}&\Delta t\\ \frac{1}{2}\Delta t^{2}&\Delta t\end{bmatrix}\emph{{P}}_{2}^{T}.$ (45)
@@
$\Omega^{T}=V_{Q}\left(t_{k},t_{k}+\Delta t/2,t_{k}+\Delta t\right)\emph{{P}}_{Q}^{T}.$ (47)
As a matter of convention, the polynomial order in our Vandermond matrix decreases with column index, such as $V_{i,j}=t_{i}^{Q-j}$. Solving for P in (45) and substituting the result in (47), we can express $\Omega$ in terms of $\Theta_{2}$
$\Omega^{T}$ $=\begin{bmatrix}0&1\\ \frac{\Delta t}{2}&1\\ \Delta t&1\\ \end{bmatrix}\begin{bmatrix}-\frac{1}{2}\Delta t^{2}&\Delta t\\ \frac{1}{2}\Delta t^{2}&\Delta t\end{bmatrix}^{-1}\Theta_{2}^{T}$ (48)
$=\frac{1}{2\Delta t}\begin{bmatrix}1&1\\ 0&2\\ -1&3\end{bmatrix}\Theta_{2}^{T}.$ (49)
Explicitly, we have
@@
Substituting the expressions from eq. (50) into the second-order RK4 approximation from eq. (40) we have
$\Delta\phi_{\textrm{RK4}}\left(t_{k}+\Delta t\right)\approx\Delta\boldsymbol{\theta}_{k}+\frac{1}{12}\left(\Delta\boldsymbol{\theta}_{k-1}\times\Delta\boldsymbol{\theta}_{k}\right).$ (51)
which is precisely the same as the classical result from eq. (30) (and derived in Appendix A)
@@
$\Theta_{3}=\begin{bmatrix}\Delta\boldsymbol{\theta}_{k-1}&\Delta\boldsymbol{\theta}_{k}&\Delta\boldsymbol{\theta}_{k+1}\end{bmatrix}$, (52)
@@
$\emph{{P}}_{3}^{T}=\begin{bmatrix}\frac{1}{3}\Delta t^{3}&-\frac{1}{2}\Delta t^{2}&\Delta t\\ \frac{1}{3}\Delta t^{3}&\frac{1}{2}\Delta t^{2}&\Delta t\\ \frac{7}{3}\Delta t^{3}&\frac{3}{2}\Delta t^{2}&\Delta t\end{bmatrix}^{-1}\Theta_{3}^{T}.$ (53)
The three samples points for $\omega$ required by the RK4 algorithms are then given by
$\Omega^{T}$ $=V_{3}\left(t_{k},t_{k}+\Delta t/2,t_{k}+\Delta t\right)\emph{{P}}_{3}^{T}$ (54a)
$=\frac{1}{24\Delta t}\begin{bmatrix}8&20&-4\\ -1&26&-1\\ -4&20&8\end{bmatrix}\Theta_{3}^{T}.$ (54b)
Replacing $\omega$ in eq. (40), using eq. (54b), we obtain
$\Delta\phi_{\textrm{RK4}}\left(t_{k}+\Delta t\right)\approx\Delta\boldsymbol{\theta}_{k}+\frac{1}{288}\left(\Delta\boldsymbol{\theta}_{k+1}\times\Delta\boldsymbol{\theta}_{k-1}+13\left(\Delta\boldsymbol{\theta}_{k-1}-\Delta\boldsymbol{\theta}_{k+1}\right)\times\Delta\boldsymbol{\theta}_{k}\right).$ (55)

### v2

@@ 3.1 Simplification of the Coning Correction
When the rotation vector is small (which it often is for small time steps since integration is always started from $\boldsymbol{\phi}=\textbf{0}$), the coefficient of the right-hand term may be expanded as a Taylor series in the same was as eq. (15) to find
@@
which is an excellent approximation in most cases. More than this, the so-called Goodman-Robinson approximation [Goodman:1958] maintains only first-order terms in the attitude and replaces $\boldsymbol{\phi}$ with $\boldsymbol{\theta}$ inside the integrand
@@
The replacement of $\boldsymbol{\phi}$ by $\boldsymbol{\theta}$ may be viewed as a single iteration within a Picard iteration scheme. The Goodman-Robinson approximation, which predates the Bortz’s work by over a decade, is the foundation for many of the classical coning algorithms [Ignagni:1990, Ignagni:1996, Savage:1998]. Some newer algorithms attempt to solve either eq. (25) or eq. (27), though this comes with an increase in complexity [Savage:2006, Wang:2019, Ignagni:2020].
@@ 3.2 Approximate Solutions to the Coning Correction Term
Consider a rate integrating gyro producing measurements of $\Delta\boldsymbol{\theta}_{k}$ at the sensor interval from eq. (22). Suppose that $m$ such measurements $\{\Delta\boldsymbol{\theta}_{k}\}_{k=1}^{m}$ are accumulated over an IMU’s minor interval. The objective is to integrate this set of IMU observations to produce an estimate of the DCM $\emph{{T}}_{m}$ or the rotation vector $\boldsymbol{\phi}_{m}$ at the end of the minor interval. The Goodman-Robinson approximation from eq. (28) may be used to solve this problem analytically if some assumptions on the time-history of $\boldsymbol{\omega}$ are made. The simplest non-trivial assumption is a linearly varying angular velocity.
Most of the classical coning literature focuses on the two-speed paradigm where $m>1$, where the “two speeds” being the rate of sensor interval and the rate of the minor interval (c.f. Fig. 1). Under the two-speed paradigm, where $m>1$ sensor interval measurements are obtained over a minor interval, a linear $\boldsymbol{\omega}$ may be assumed to obtain the classic approximation [Ignagni:1990]
@@
The two-speed techniques exemplified by eq. (29)—though there are many others under different assumptions for the time-history of $\boldsymbol{\omega}$ [Ignagni:1996, Savage:1998]—are exceptionally fast and remove the need for performing the computationally expensive task of finite rotations at the sensor interval. However, there have been considerable advancements in computing since 1990, and two-speed algorithms are not always necessary. Propagating at the sensor interval is well-known to be desirable when it is possible [Savage:1998].
If finite rotations are to be performed at the sensor interval, then coning correction is only applied over the sensor interval (and not the minor interval). This leads to the one-speed algorithm, which may be easily derived for linear $\boldsymbol{\omega}$ as described in the Appendix and results in
@@
which is the same correction suggested by Miller in the early 1980s [Miller:1983]. Thus, letting $\Delta\emph{{T}}_{k}=\emph{{T}}(\Delta\boldsymbol{\phi}_{k})$ be computed from eq. (4), the attitude at the end of the sensor interval is then computed as the product
$\emph{{T}}_{m}=\prod_{k=1}^{m}\Delta\emph{{T}}_{k}=\Delta\emph{{T}}_{m}\cdots\Delta\emph{{T}}_{k}\cdots\Delta\emph{{T}}_{1}$ (32)
@@ 4 Numerical Methods for Coning Correction
In light of eq. (20), coning compensation amounts to solving the differential equation of eq. (13) for $\boldsymbol{\phi}\left(t_{k}+\Delta t\right)$ from an initial condition $\boldsymbol{\phi}\left(t_{k}\right)=0$. This generalized perspective enables usage of traditional numerical solvers such as Runge-Kutta methods. Indeed, using traditional Runge-Kutta style solvers an identical expression is obatined for the single-speed algorithm fom eq. (31). There are three benefits to exploring an alternative derivation. First, equivalence with traditional coning correction methods demonstrates applicability of Runge-Kutta style solvers. Second, generalized solvers accommodate non-zero initial conditions. Solving for $\Delta\boldsymbol{\phi}_{k}$ only requires solving the Bortz equation from an initial condition of all zeros. A generalized approach facilitates solving for $\boldsymbol{\phi}_{k}$ from $\boldsymbol{\phi}_{k-1}$ directly potentially reducing computations. Thirdly, the recursive structure of Runge-Kutta methods easily generalizes to higher-order algorithms.
@@ 4.1 Runge-Kutta Methods and Sampling Requirements
The Runge-Kutta algorithm, stated in Algorithm 1, is a generalized numerical solver for initial value problems. The generalized ordinary differential equation $f(t,\boldsymbol{\phi}):\mathbb{R}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, is solved from an initial condition ${\boldsymbol{\phi}}_{k}\in\mathbb{R}^{3}$ at time $t_{k}$ to the final value ${\boldsymbol{\phi}}_{k+1}={\boldsymbol{\phi}}\left(t_{k}+\Delta t\right)$ at stepsize $\Delta t$.
@@
Algorithm 1 extends to different solver orders and sampling schemas as determined by $\emph{{A}}\in\mathbb{R}^{n\times n}$, $\emph{{b}},\emph{{c}}\in\mathbb{R}^{n}$. Butcher tableaux [butcherRungeKuttaProcessesHigh1964, butcherNumericalMethodsOrdinary2016] are useful for depicting the algorithm hyperparameters A,b,c with examples shown in Figs. 3 - 6 for common implementations. Note, the array c, depicted in tableau’s first column, identifies both the solver order and the timesteps at which point the differential equation, $f$, is evaluated. The sample points are listed relative to step size $\Delta t$ (c.f. $t_{0}+\Delta t\emph{{c}}_{\nu}$} in Algorithm 1, line 5).
@@
Figure 3: Forward Euler Tableau
@@
Figure 4: Explicit-Midpoint Tableau
@@
Figure 5: Runge-Kutta Original Third Order Tableau
@@
Figure 6: Runge-Kutta Original Fourth Order Tableau
To solve eq. (20), using Algorithm 1, eq. (13) is employed as
$f(t,\boldsymbol{\phi})\coloneq\emph{{J}}^{-1}_{\boldsymbol{\phi}}\boldsymbol{\omega}(t)$ (33)
In this case, the $f$ separates as two terms. The first term, $\emph{{J}}^{-1}_{\boldsymbol{\phi}}$, depends on the dependent variable $\boldsymbol{\phi}$. The second term $\boldsymbol{\omega}(t)$, for inertial navigation, is obtained through sensor measurements. Notice, Algorithm 1 samples $\boldsymbol{\omega}(t)$ at discrete points as determined by $t_{k}$, $\Delta t$, and c. First Runge-Kutta methods are expanded assuming direct access to $\boldsymbol{\omega}$ for any $t$.
Applying the forward-Euler method, calling Algorithm 1 using Fig. 3, yields
@@
Motivated by eq. (20), $\boldsymbol{\phi}_{k}=0$ is selected which effects $\emph{{J}}_{\boldsymbol{\phi}_{k}}^{-1}=\emph{{I}}_{3\times 3}$}. Consequently,
@@
For higher-order methods, expanding Algorithm (1) becomes tedious. However, using the initial condition $\boldsymbol{\phi}(t_{k})={\bf 0}$, the small-angle approximation eq. (16), and retaining only quadratic terms,
@@
are obtained. Samples of $\boldsymbol{\omega}$ must be provided either through direct measurements (e.g., gyros providing instantaneous rates) or from a model (e.g., parameterized from integrated rate measurements).
@@ 4.2 Approximating $\boldsymbol{\omega}$ from $\Delta\boldsymbol{\theta}$
When $\boldsymbol{\omega}$ is available, eq. (41) may be used directly. However, for rate-integrating gyros, $\boldsymbol{\omega}$ must be modeled from $\Delta\boldsymbol{\theta}$ measurements. Consider a model for $\boldsymbol{\omega}$ that is polynomial in the elapsed time
$\boldsymbol{\omega}\left(t\right)=\sum_{i=1}^{Q}\emph{{p}}_{i}\left(t-t_{k}\right)^{i-1}$ (42)
The model parameters $P_{Q}\in\mathbb{R}^{3\times Q}$
$P_{Q}=\begin{bmatrix}\emph{{p}}_{1}&\emph{{p}}_{2}&\cdots&\emph{{p}}_{Q}\end{bmatrix}$ (43)
can be obtained from $\Delta\boldsymbol{\theta}$ measurements by solving a linear system of equations. For each measurement, equation by is obtained by replacing $\boldsymbol{\omega}$ in eq. (22) with eq. (42)
@@
$=\sum_{i=0}^{Q}\emph{{p}}_{k}\frac{1}{i}\left(\left(j\Delta t\right)^{i}-\left(j-1\right)^{i}\Delta t^{i}\right)$ (44b)
From two measurements, an affine model can be formed for $\boldsymbol{\omega}$. Collecting two measurements in the matrix $\Theta_{2}\in\mathbb{R}^{3\times 2}$
$\Theta_{2}=\begin{bmatrix}\Delta\boldsymbol{\theta}_{k-1}&\Delta\boldsymbol{\theta}_{k}\end{bmatrix}$ (45)
a linear system of equations is formed
$\Theta^{\top}=\begin{bmatrix}-\frac{1}{2}\Delta t^{2}&\Delta t\\ \frac{1}{2}\Delta t^{2}&\Delta t\end{bmatrix}P_{2}^{\top}$ (46)
@@
$\Omega^{\top}=V_{Q}\left(t_{k},t_{k}+\Delta t/2,t_{k}+\Delta t\right)P_{Q}^{\top}$ (48)
As a matter of convention, the polynomial order in the Vandermonde matrix decreases with column index, such that $V_{i,j}=t_{i}^{Q-j}$. Solving for $P$ in (46) and substituting the result in (48), $\Omega$ is expressed in terms of $\Theta_{2}$.
$\Omega^{\top}$ $=\begin{bmatrix}0&1\\ \frac{\Delta t}{2}&1\\ \Delta t&1\\ \end{bmatrix}\begin{bmatrix}-\frac{1}{2}\Delta t^{2}&\Delta t\\ \frac{1}{2}\Delta t^{2}&\Delta t\end{bmatrix}^{-1}\Theta_{2}^{\top}$ (49)
$=\frac{1}{2\Delta t}\begin{bmatrix}1&1\\ 0&2\\ -1&3\end{bmatrix}\Theta_{2}^{\top}$ (50)
Explicitly,
@@
Substituting the expressions from eq. (51) into the second-order RK4 approximation from eq. (41) reveals
$\Delta\boldsymbol{\phi}_{\textrm{RK4}}\left(t_{k}+\Delta t\right)\approx\Delta\boldsymbol{\theta}_{k}+\frac{1}{12}\left(\Delta\boldsymbol{\theta}_{k-1}\times\Delta\boldsymbol{\theta}_{k}\right)$ (52)
which is precisely the same as the classical result from eq. (31) (and derived in the Appendix).
@@
$\Theta_{3}=\begin{bmatrix}\Delta\boldsymbol{\theta}_{k-1}&\Delta\boldsymbol{\theta}_{k}&\Delta\boldsymbol{\theta}_{k+1}\end{bmatrix}$ (53)
@@
$P_{3}^{\top}=\begin{bmatrix}\frac{1}{3}\Delta t^{3}&-\frac{1}{2}\Delta t^{2}&\Delta t\\ \frac{1}{3}\Delta t^{3}&\frac{1}{2}\Delta t^{2}&\Delta t\\ \frac{7}{3}\Delta t^{3}&\frac{3}{2}\Delta t^{2}&\Delta t\end{bmatrix}^{-1}\Theta_{3}^{\top}$ (54)
The three samples points for $\boldsymbol{\omega}$ required by the RK4 algorithms are then given by
$\Omega^{\top}$ $=V_{3}\left(t_{k},t_{k}+\Delta t/2,t_{k}+\Delta t\right)P_{3}^{\top}$ (55a)
$=\frac{1}{24\Delta t}\begin{bmatrix}8&20&-4\\ -1&26&-1\\ -4&20&8\end{bmatrix}\Theta_{3}^{\top}$ (55b)
Replacing bomegabomega in eq. (41), using eq. (55b), results in
$\Delta\boldsymbol{\phi}_{\textrm{RK4}}\left(t_{k}+\Delta t\right)\approx\Delta\boldsymbol{\theta}_{k}+\frac{1}{288}\left(\Delta\boldsymbol{\theta}_{k+1}\times\Delta\boldsymbol{\theta}_{k-1}+13\left(\Delta\boldsymbol{\theta}_{k-1}-\Delta\boldsymbol{\theta}_{k+1}\right)\times\Delta\boldsymbol{\theta}_{k}\right)$ (56)
The expression (56) is an approximation to RK4 integration where the required samples of $\boldsymbol{\omega}$ were obtained by fitting a curve to three samples of $\Delta\boldsymbol{\theta}$. This approach could be extended in several ways. For example, additional samples such as $\Delta\boldsymbol{\theta}_{k-2}$ could yield higher-order curve fit for $\boldsymbol{\omega}$. Alternatively, curve fitting could be applied in a least-squares sense to mitigate noise or quantization effects in $\Delta\boldsymbol{\theta}$ measurements.